What is the standard form of a quadratic equation for the following problem and how can I understand it? (48n +3)(n-2) -48 = 22
Answer by stanbon(75887)

What is the standard form of a quadratic equation for the following problem and how can I understand it?
(48n +3)(n-2) -48 = 22
-------------------
48n^2-96n+3n-6 - 48 = 22
-----
48n^2-93n = 22+54
---
48n^2 - 93n - 76 = 0
---
I graphed the quadratic and found
solutions at x = 2.5568 and x = -0.6193
---
You could also use the Quadratic Formula to solve.
===========
Cheers,
Stan H.